OpenStudy (anonymous):

Rephrasing a question! Can someone tell me where I'm going wrong? I'm supposed to factor this: −3z2+5z+2 And get this: -3(z+1/3)(z-2) This is what I've got so far: −3z2+5z+2 3(z2+5z/3+2/3) −(5z/3)/2+−(5z/3)/2−2/3−−−−−−−−−−−−√ x1=2/3−−−√ Aaannd here it breaks down...

OpenStudy (anonymous):

This step: −(5z/3)/2+−(5z/3)/2−2/3−−−−−−−−−−−−√ Is supposed to read: \[-(5z/3)/2+-\sqrt{(5z/3)/2-2/3}\]

OpenStudy (anonymous):

Nope, it's +2 at the end. :)

OpenStudy (anonymous):

I realize there may be several correct answers, but what my book wants and I don't get how to arrive at is -3(z+1/3)(z-2)

OpenStudy (anonymous):

Nope, it only gives a hint. Start by breaking out -3. So the first step seems to be -3(z2+5z/3+2/3)

OpenStudy (anonymous):

polynomials in factored form

OpenStudy (anonymous):

I've only got one identical problem. This: \[7x^2 - 5x -2\] Apparently (?) factors into this: \[7(x - 1)(x + 2/7)\]

OpenStudy (anonymous):

I think part of the solution is the application of the following formula: \[x = -p/2 +- \sqrt{(p/2)^2-q}\]

OpenStudy (anonymous):

But, I don't know, haven't solved it...

OpenStudy (anonymous):

This is actually fun, I just wish I didn't get stuck all the time!

OpenStudy (anonymous):

Hmm, I get this: \[7x^2-5x-2\] \[-(5/7)/2 +- \sqrt{((5/7)/2)^2+2/7}\] \[x1: \sqrt{2/7}\] And stuck... the answers should be 2/7 and -1

OpenStudy (anonymous):

well yeah, but x1 is already wrong. I should be getting 2/7, not sqrt(2/7)

OpenStudy (anonymous):

There's a facepalm moment coming up soon, I can feel it.

OpenStudy (anonymous):

I can't think of a way of factoring x^2 -5x/7-2/7 without it

OpenStudy (anonymous):

thanks, I'll watch it

OpenStudy (jiteshmeghwal9):

\[\Huge{-3z^2-6z+z+2}\]now take common from \(\Large{(-3z^2-6z)}\) & \(\Large{(z+2)}\) i.e.;\[\Huge{-3z(z+2)+1(z+2)}\]now take common \[\Huge{(-3z+1)(z+2)Ans.}\]

OpenStudy (jiteshmeghwal9):

gt it @math0101 ???

OpenStudy (jiteshmeghwal9):

@math0101 the answer i have founded is correct . you can check http://www.wolframalpha.com/input/?i=factor+%3A+%28-3z^2%2B5z%2B2%29

OpenStudy (jiteshmeghwal9):

click on this highlighted site :)

OpenStudy (anonymous):

that may be, and I appreciate it, but I still don't know how to arrive at the answer the book wants :/

OpenStudy (anonymous):

interesting site btw

OpenStudy (hba):

Do u want me to explain it to you

OpenStudy (jiteshmeghwal9):

yeah ! computing site :

OpenStudy (jiteshmeghwal9):

can u do it @hba ???

OpenStudy (anonymous):

hba: if you please ;)

OpenStudy (hba):

Yeah it easy ill do it wait

OpenStudy (hba):



OpenStudy (hba):